OpenStudy (anonymous):

What do the following two equations represent? −5x+5y=−1 25x−25y=5 perpendicular parallel equivalent none of the above

OpenStudy (anonymous):

plug the 2 equations into Wolfram alpha or graph them

OpenStudy (anonymous):

What do the following two equations represent? −x+3y=3 −3x−y=2

OpenStudy (anonymous):

what about this one

OpenStudy (anonymous):

your going to have to graph

OpenStudy (anonymous):

or find the slope of each